At school, 40% of students drink coffee. Consider a random sample of 50 students. What is the pro...

  • At school, 40% of students drink coffee. Consider a random sample of 50 students. What is the probability that less than 30% of these students drink coffee?

(Show all applicable work)
Answer #1

(0)= 0、40 1 So n: 50 so mean (4), np = 50×0.40 =0-60 20 p(2 L-1.443)